The present value of a loan that calls for the payment of 5,000 per year for 6 years if the discount rate is 10 percent and the first payment will be made one year from now? How many would your answer change if the 5000 per year occurred for 10 years
Answer by Theo(13342)

the formula you need to use is:

PRESENT VALUE OF AN ANNUITY WITH END OF TIME PERIOD PAYMENTS

p = (a*(1-1/(1+r)^n))/r

p is the present value of the annuity.
a is the annuity.
r is the interest rate per time period.
n is the number of time periods.

a = 5000
r = .10 per year
n = first 6 years and then 10

with 6 years, the formula becomes:
p = (5000 * (1 - 1/(1.10)^6))/.10
solve for p to get p = 21776.3035

with 10 years, the formula becomes:
p = (5000 * (1 - 1/(1.10)^10))/.10
solve for p to get p = 30722.83553

i confirmed using a financial calculator that these answers are correct.

when you are using these formulas in your calculator, you need to use all the parentheses exactly the way they are shown.